OpenStudy (anonymous):

if x men can do a job in h days, ow long would y men take to do the same job?

OpenStudy (anonymous):

\[A) \frac{ x}{ h }\] \[B) \frac{ xh}{ y }\] \[C) \frac{hy }{ x }\] \[D) \frac{ xy}{ h }\] \[E) \frac{ x}{y }\]

OpenStudy (anonymous):

\[j=\frac{ hy }{ x }?\]

OpenStudy (mathmate):

x men can do a job in h days, so 1 man will do the job in h*x days. How many days would y men take to do the job?

OpenStudy (anonymous):

i think its letter C

OpenStudy (mathmate):

Explain your reasoning please.

OpenStudy (anonymous):

well i used what hero did and got \[\frac{ jx }{ y }=h \rightarrow (y)\frac{ jx }{ y }=h \]

OpenStudy (anonymous):

oh.....

OpenStudy (anonymous):

so its not\[j=\frac{ hy }{ x }\]?

OpenStudy (anonymous):

ok

OpenStudy (mathmate):

Try to follow this reasoning: x men can do a job in h days, so 1 man will do the job in h*x days. (more men=faster, less men=slower). How many days would y men take to do the job? Can you figure it out?

OpenStudy (anonymous):

not really...

OpenStudy (mathmate):

We'll look at it this way: x men can do a job in h days. The job requires x*h man-days to finish. How many days would it take y men to do a job requiring x*h man-days?

OpenStudy (mathmate):

@ineedhelpnow08 are you still there?

OpenStudy (anonymous):

yeah....just thinking

OpenStudy (mathmate):

The idea of man-days makes the calculation easy. A job requiring 30 man-days can be done by 5 men in \(\large \frac{30 man-days}{5 men}=6 days \). Notice how the numbers AND the units cancel. Similarly, a 30 man-day job can be done by 2 men in 30/2=15 days.

OpenStudy (anonymous):

hmmm so like if i use this \[y \times k=x \times h\] would it work?

OpenStudy (mathmate):

Yes, that's the idea. What do you get?

OpenStudy (anonymous):

\[\frac{ xh }{ y }\]

OpenStudy (mathmate):

Are you sure?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

i divided y on each side to find k

OpenStudy (mathmate):

Great! Just checking...! congrats.

OpenStudy (anonymous):

thank you!!!

OpenStudy (mathmate):

You're welcome!

OpenStudy (anonymous):

life saver :)